# IndoCulture: Exploring Geographically-Influenced Cultural Commonsense Reasoning Across Eleven Indonesian Provinces

### v1

@@ 3.2 Data Statistics
After the annotation process, we initially collected 3,162 instances. [^4] Although we requested each annotator to produce 150 instances, not all were able to complete their allotted tasks within the given timeframe. Unfortunately, we were unable to find additional candidates from the same local province to address the data deficiencies Winata et al. (2023). In stage 1 of quality control, the initial pool of 3,162 instances was reduced to 2,801 instances. Our thorough stage 2 of quality control further reduced the sample to 2,429 final high-quality samples. The data distribution of IndoCulture per province is presented in Table 1. Approximately three-quarters of IndoCulture instances contain province-specific content, with an average length of around 35 words. IndoCulture covers multiple topics, as illustrated in Figure 2.
@@ 4.1 Set-Up
We focus on evaluating 23 language models in zero-shot settings: (1) 15 open-source multilingual language models including BLOOMZ Muennighoff et al. (2022), mT0 Muennighoff et al. (2022), LlaMA2 Touvron et al. (2023), Bactrian-X Li et al. (2023a) in different sizes, (2) 2 South East Asian language models including SeaLLM Nguyen et al. (2023b), and SeaLion Singapore (2023), (3) 4 Indonesian-centric language models, including IndoBART Cahyawijaya et al. (2021), IndoGPT Cahyawijaya et al. (2021), Merak Ichsan (2023), Komodo Owen et al. (2024), (3) 2 closed-source models, including ChatGPT: gpt-3.5-turbo Ouyang et al. (2022) and GPT–4: gpt-4-0613 OpenAI (2023). Please refer to Appendix A for further details.
Initially, we evaluate the effectiveness of sentence completion and multiple-choice question strategies in predicting the correct options. In both scenarios, we conduct benchmarks across three distinct location contexts. Formally, given a premise $s$, three candidate options $c_{1}$, $c_{2}$, $c_{3}$, and location $l\in\{\texttt{none},\texttt{Indonesia},p_{j}\}$, for sentence completion, we select the correct option based on:
@@
Here, $\text{concat}(s,c)$ denotes the concatenation of premise $s$ and candidate option $c$, separated by a space. In the case of multiple-choice questions, we devise a template for the prompt question and determine the answer by selecting the option with the highest probability among letters A, B, and C. Figure 3 displays the prompt templates in Indonesian and English for both sentence completion and multiple-choice question.
@@ 4.2 Results
BLOOMZ (560M) 37.5 35.3 35.3 32.1 32.0 32.1
BLOOMZ (1.1B) 36.2 36.9 37.2 32.0 32.0 32.0
BLOOMZ (3B) 38.5 40.5 41.2 45.9 47.6 48.6
BLOOMZ (7B) 40.9 43.8 44.4 48.8 49.8 49.7
@@
LLaMA2 (7B) 36.7 37.0 37.2 40.0 39.0 37.7
LLaMA2-chat (7B) 36.9 36.9 37.4 40.5 41.1 40.3
LLaMA2 (13B) 39.4 39.7 39.9 46.9 46.9 46.6
LLaMA2-chat (13B) 38.2 38.4 38.9 47.2 49.2 49.2
@@
IndoBART (132M) 41.9 41.1 41.7 32.1 32.0 32.2
IndoGPT (117M) 42.6 42.1 42.7 33.7 34.1 35.0
Merak (7B) 40.2 40.6 42.6 51.5 52.5 52.5
SeaLLM (7B) 39.1 39.7 41.1 51.7 52.4 52.5
SEA-LION (7B) 38.8 38.7 39.6 33.4 32.8 33.0
@@
GPT–3.5 (NA) 58.4 60.2 61.7
GPT–4 (NA) 68.8 71.5 75.8
Table 2: Zero-shot accuracy across various models and settings. “MCQ” refers to the multiple-choice question method, and $l$ denotes the location as additional context. The bold numbers highlight the highest score within each model group.
@@
Multiple-choice question method is generally better.
Our findings suggest that the multiple-choice question method tends to outperform the sentence completion method, with exceptions noted for BLOOMZ (560M, 1.1B), IndoBART, IndoGPT, and Komodo. Interestingly, within the sentence completion method, the small Indonesian-centric language model, IndoGPT, surpasses the performance of almost all large multilingual models and remains competitive with BLOOMZ (7B). However, IndoGPT’s performance experiences a significant decline in the multiple-choice question method, with a notable margin of nearly 8 to 10 points. While Komodo (7B) achieves the highest accuracy with the sentence completion method, its performance also decreases by 10-12 points in the multiple-choice method. This discrepancy could potentially be attributed to differences in the nature of language model training and instruction-tuning.
@@
Our investigation reveals that incorporating various levels of location granularity has a noticeable effect on zero-shot performances, especially models with larger parameter sizes. Detailed location contexts notably enhance the accuracy of BLOOMZ (7B), LlaMA2 (13B), Merak (7B), SeaLLM (7B), Komodo (7B), GPT–3.5, and GPT–4. For instance, in GPT–4, the accuracy gap between $l=\texttt{none}$ and $l=\texttt{Indonesia}$ is 2.7, and this gap further increases by 7 points when $l=\texttt{Prov}$ is assigned.
@@ 4.3 Analysis
Given the superior performance of models with large parameter sizes using the multiple-choice question method and location $l=\texttt{Prov}$, we employ these settings for analysis.
Province Merak GPT–3.5 GPT–4
@@
Aceh 59.7 53.4 79.2 65.5 93.1 73.6
North Sumatra 58.3 46.9 63.9 57.2 75.0 73.2
West Sumatra 50.0 41.3 63.2 48.4 85.5 63.7
West Java 65.3 58.8 84.2 71.8 92.6 81.7
Cental Java 58.8 47.1 67.6 52.9 82.4 72.9
East Java 57.4 37.6 72.2 46.5 87.0 63.4
Bali 78.9 65.6 87.7 81.7 93.0 89.4
NTT 64.3 52.0 57.1 49.3 85.7 68.0
South Borneo 65.7 46.4 68.6 53.6 77.1 67.4
South Sulawesi 53.3 47.9 66.7 55.8 80.0 70.6
Papua 76.7 55.2 76.7 62.8 90.0 71.3
Table 3: Top-3 model accuracy by province. "PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
@@
Table 3 highlights that the top 3 performing LLMs exhibit a nuanced understanding of cultures within Indonesian provinces, particularly excelling in the cultures of West Java and Bali compared to other provinces. GPT–4, for instance, achieves the best accuracy of 93%, while Merak and GPT–3.5 also demonstrate strong performance in Bali and West Java, with accuracies ranging between 55% and 88%. In other provinces like West Sumatra and South Borneo, the models typically exhibit poorer performance. Specifically, for GPT-3.5, the performance gap compared to Bali ranges from 19 to 33%. This highlights the presence of cultural biases and a lack of inclusivity in model reasoning abilities, likely stemming from the distribution of training data. The proximity of West Java to Jakarta (Indonesia’s capital) and Bali’s global status as a tourism destination may contribute to the abundance of textual data on these two cultures.
@@
Table 4 shows the accuracy of the top 3 performing models across topics. Similar to Table 3, the models perform better in province-specific context for all topics, with death customs being an exception. For province-specific contexts, GPT–4 excels in themes of wedding, death, and arts while for non-specific contexts, GPT–4 achieves accuracy more than 90.0 for family relationship, religious holidays, and farm and fishery.
Province Merak GPT–3.5 GPT–4
@@
Food 56.9 51.0 65.9 62.1 79.3 63.5
Wedding 66.7 64.3 77.6 56.0 92.8 78.6
Family relationship 65.5 54.5 72.5 53.9 93.1 72.7
Pregnancy and kids 67.9 42.9 69.2 50.6 82.1 72.7
Death 38.5 55.2 60.9 62.5 76.9 82.5
Religious holidays 68.5 48.8 79.3 68.8 94.4 64.6
Agriculture 70.5 52.3 71.4 62.1 88.5 76.1
Farm and fishery 77.8 51.0 77.8 51.9 94.4 75.8
Art 58.7 50.2 85.2 61.1 78.3 77.3
Traditional game 60.0 50.2 80.3 59.8 88.0 66.9
Daily activities 55.1 39.6 66.0 64.8 87.8 63.0
Socio-religious 57.1 44.3 79.6 61.1 85.7 64.6
Table 4: Top-3 model accuracy by topic. "PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
@@
We automatically translated IndoCulture to English using the Google Translate API [^5] Accessed in March 2024. and used the English prompt in Figure 3 to evaluate the models. All model performances in English text dropped except for LLaMA-2 and Merak. This could be attributed to two reasons. First, LLaMA-2 is an English-centric model, and Merak is pretrained from LLaMA-2. Second, the performance drop in other models could be caused by translation errors. We further investigated this with 100 random samples and found that 81 samples had acceptable translations. We observed translation errors such as pronoun mismatches, inaccurate proverb translations, and inaccurate translations of local terms, such as “pupuik” translated as “fertilizer”.666pupuik is a traditional musical instrument in West Sumatra. It is worth noting that the word pupuik closely resembles the word pupuk in Indonesian, which means fertilizer. To better understand the cultural gap in language models, we followed the approach of Liu et al. (2023a) to manually correct the translations and reevaluate the models. We found that GPT-4’s performance over the 100 random samples was 77.0 for the original Indonesian text, 68.0 for the English machine-translated text, and 72.0 for the English translation fixed by humans.
@@
Results by fine-grained culture elements
We tasked two expert workers with annotating 200 random samples from IndoCulture based on six cultural elements, as derived from Axtell and Fornwald (1998); Williams (2014). [^7] https://www.languageeducatorsassemble.com/elements-of-culture [^8] https://pressbooks.howardcc.edu/soci101/chapter/3-2-the-elements-of-culture/ While these elements may not encompass every cultural aspect, we contend that they cover the most prominent or pivotal elements, including: (1) symbols (material or non-material objects representing meaning), (2) artifacts (material or non-material objects produced by society), (3) values and beliefs (principles, ideas, and concepts assumed to be ideal and correct in society), (4) norms (rules guiding values and beliefs), (5) language, (6) rituals (established procedures and ceremonies), and (7) other, for examples that do not fit into any of the defined elements. This annotation is a multi-label task, and the average Kappa score across the cultural elements is 0.56, with each ranging from 0.4 to 0.75. These scores indicate moderate to substantial agreement.
Table 5 displays the distribution of each cultural element in our dataset, along with the performance breakdown across Merak, GPT-3.5, and GPT-4. Among the 200 random samples, we observe that 42.5% of our data contains artifacts, 37.5% norms, and 30% rituals. Only 4% of the data pertains to symbols, while 7.5% belongs to the other category. Merak exhibits lower performance in norms, with a 24% decrease compared to values and beliefs. Conversely, GPT-3.5 and GPT-4 perform better in norms, with the performance gap in values and beliefs being only $-4$ points. Additionally, language poses a challenge in Merak, with only 38% accuracy, while GPT-3.5 and GPT-4 achieve 60% and 72% accuracy, respectively.
Cultural element (%) Merak GPT–3.5 GPT-4
Symbols (4) 50.0 58.9 70.8
Artefacts (42.5) 55.3 59.1 67.8
Values and Beliefs (10.5) 61.9 57.5 69.8
Norms (37.5) 38.7 63.2 73.6
Language (19.5) 38.5 60.2 72.0
Ritual (30) 53.3 60.0 70.7
Other (7.5) 66.7 58.9 69.2
Table 5: Accuracy comparison of Merak, GPT–3.5, and GPT–4 across 200 random samples, categorized by cultural elements. The numerical value following each cultural element indicates its proportion within the samples.
Can the model provide a reasonable explanation to support the answer?
Figure 5: Performance comparison between Merak (7B) and GPT–4 based on text generation output. “Answer (T)” indicates that the generated answer is true, while ”Exp(F)” denotes that the answer explanation is false.
We conduct a manual investigation of the text generation output for Merak (7B) and GPT–4 across 200 random samples. This involves manually examining the generated answer along with its explanation. To obtain explanation, we modify the Indonesian prompt in Figure 3 by adding the string Jelaskan jawabanmu! ("Explain your answer!"). Our annotation process is binary, categorizing explanations as either True or False. We label an explanation as False if it is absent, contains hallucinations, or provides inaccurate information. [^9] We use Google search engine to verify the correctness of the explanation.
As anticipated, there is a significant drop in accuracy for Merak (7B) from 52.5% (as shown in Table 2) to 29.5%. This discrepancy underscores the limitations of relying solely on token probability to assess the true capability of a language model. Interestingly, only 4.5% of the samples are answered correctly with the appropriate explanation by Merak, despite it being the top performer among the open-source models. GPT–4 achieves an accuracy of 69.5%, which is 5 points lower than indicated in Table 2, possibly due to a smaller sample size and biases in regular expressions. However, GPT–4 provides better explanations, with 51.5% of the samples correctly answered along with the appropriate explanation.

### v2

@@ 3.2 Data Statistics
After the instance writing process, we initially collected 3,162 instances out of a target of 3,300 instances (22 workers $\times$ 150 subtopics). Although we requested each annotator to produce 150 instances, not all were able to complete their allotted tasks within the given timeframe. Unfortunately, we were unable to find additional candidates from the same local province to address the data deficiencies Winata et al. (2023).
In stage 1 of quality control, the initial pool of 3,162 instances was reduced to 2,801 instances, and stage 2 of quality control further reduced the sample to 2,429 high-quality samples. The data distribution of IndoCulture per province is presented in Table 2. Approximately three-quarters of IndoCulture instances contain province-specific content, with an average length of around 35 words. IndoCulture covers multiple topics, as illustrated in Figure 2.
@@ 4.1 Set-Up
We evaluate 27 language models in zero-shot settings: (1) nineteen open-weight multilingual language models of varying sizes, namely BLOOMZ Muennighoff et al. (2023), mT0 Muennighoff et al. (2023), Bactrian-X Li et al. (2023), Llama–2 Touvron et al. (2023), and Llama–3 Dubey et al. (2024); (2) two South East Asian language models, namely SeaLLM Nguyen et al. (2024), and SeaLion Singapore (2023); (3) four Indonesian-centric language models, namely IndoBART Cahyawijaya et al. (2021), IndoGPT Cahyawijaya et al. (2021), Merak Ichsan (2023), and Komodo Owen et al. (2024); (3) two closed-weight models, namely ChatGPT: gpt-3.5-turbo Ouyang et al. (2022) and GPT–4: gpt-4-0613 OpenAI (2023). Please refer to Appendix A for further details.
First, we evaluate the effectiveness of sentence completion and multiple-choice question strategies in predicting the correct options using the Indonesian and English prompt templates shown in Figure 3. In both scenarios, we conduct benchmarks across three distinct location contexts. Formally, given a premise $s$, three candidate options $c_{1}$, $c_{2}$, $c_{3}$, and location $l\in\{\texttt{none},\texttt{Indonesia},province\}$, for sentence completion, we select the correct option based on:
@@
Here, $\text{concat}(s,c)$ denotes the concatenation of premise $s$ and candidate option $c$, separated by a space. In the case of multiple-choice questions, we devise a template for the prompt question and determine the answer by selecting the option with the highest probability among letters A, B, and C.
For GPT–3.5 and GPT–4, we exclude experiments with sentence completion because the closed-weight models do not provide an overall probability score. For multiple-choice questions, we select the first generated token that corresponds to the letters A, B, or C using a regular expression.
@@ 4.2 Results
BLOOMZ (560M) 37.2 35.3 35.3 32.5 32.4 32.5
BLOOMZ (1.1B) 36.3 36.9 37.2 32.4 32.4 32.4
BLOOMZ (3B) 38.6 40.7 41.5 47.0 48.6 49.2
BLOOMZ (7B) 41.3 44.1 44.6 49.5 50.6 50.5
@@
[…]
Llama–2 chat (7B) 37.3 37.4 37.9 40.6 41.3 40.7
Llama–2 (13B) 39.6 40.2 40.2 47.6 47.6 47.3
Llama–2 chat (13B) 38.6 38.9 39.3 47.8 49.6 49.6
[…]
Llama–3 Instruct (8B) 41.9 41.5 42.3 56.7 57.6 59.0
Llama–3 (70B) 51.2 51.7 54.3 68.6 69.9 72.7
Llama–3 Instruct (70B) 49.2 49.6 52.2 68.5 69.3 73.3
IndoBART (132M) 42.4 41.3 42.1 32.6 32.4 32.7
IndoGPT (117M) 42.6 41.9 42.4 33.7 33.8 34.7
Merak (7B) 41.0 41.5 43.5 51.9 53.1 53.2
SeaLLM (7B) 39.1 39.3 41.1 52.2 53.1 53.0
SEA-LION (7B) 38.8 38.9 39.7 33.8 33.0 33.3
@@
GPT–3.5 (NA) 59.8 60.9 62.7
GPT–4 (NA) 69.1 71.8 75.9
Table 3: Zero-shot accuracy across various models and settings. “MCQ” refers to the multiple-choice question method, and $l$ denotes the location as additional context (“Ind” and “Prov” denote the country of Indonesia, and the corresponding province). The bold numbers highlight the highest score within each model group.
@@
The multiple-choice question method is generally better.
Our findings suggest that the multiple-choice question method tends to outperform the sentence completion method, with exceptions noted for BLOOMZ (560M, 1.1B), IndoBART, IndoGPT, and Komodo. Interestingly, in the sentence completion task, the Indonesian-focused language model Komodo (7B) outperforms nearly all large multilingual models, with the exception of Llama–3 (70B). However, Komodo experiences a significant decline for the multiple-choice question method, with a notable margin of 10–12 points. This discrepancy could potentially be attributed to differences in the nature of language model training and instruction-tuning.
@@
Our investigation reveals that incorporating various levels of location granularity has a noticeable effect on zero-shot performance, especially for models with larger parameter sizes. Detailed location context notably enhances the accuracy of BLOOMZ (7B), Llama–2 (13B), Llama–3 (70B), Merak (7B), SeaLLM (7B), Komodo (7B), GPT–3.5, and GPT–4. For instance, in GPT–4, the accuracy gap between $l=\texttt{none}$ and $l=\texttt{Indonesia}$ is 2.7, and this gap further increases by 7 points when $l=\texttt{Province}$ is assigned.
@@ 4.3 Analysis
Given the exceptional performance of models with large parameter sizes using the multiple-choice question method and location $l=\texttt{Province}$, we employ these configurations for our analysis. In this section, our main focus is on the top three performing models: Llama–3 Instruct (70B), Merak (7B), and GPT–4. These models represent a multilingual open-weight model, Indonesian-centric open-weight model, and closed-weight model, respectively.
Province Merak Llama–3 GPT–4
@@
Aceh 59.7 53.4 77.8 68.4 93.1 73.6
North Sumatra 58.3 46.9 69.4 67.5 75.0 73.2
West Sumatra 50.0 41.3 73.7 64.6 85.5 63.7
West Java 65.3 58.8 91.6 86.3 92.6 81.7
Cental Java 58.8 47.1 79.4 67.1 82.4 72.9
East Java 57.4 37.6 88.9 57.4 87.0 63.4
Bali 78.9 65.6 96.5 86.7 93.0 89.4
NTT 64.3 52.0 78.6 61.3 85.7 68.0
South Borneo 65.7 46.4 85.7 66.9 77.1 67.4
South Sulawesi 53.3 47.9 73.3 66.9 80.0 70.6
Papua 76.7 55.2 90.0 73.1 90.0 71.3
Table 4: Top-3 model accuracy by province. “PS” indicates instances containing province-specific context, while “$\neg$¬PS” indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
@@
Table 4 highlights that the top 3 performing LLMs exhibit a nuanced understanding of culture within Indonesian provinces, particularly excelling in the cultures of West Java and Bali compared to other provinces. Llama–3 and GPT–4, for instance, achieve the best accuracy of more than 90%, while Merak achieves the best performance in Bali, Papua, and West Java, with accuracies ranging between 55% and 79%. In other provinces like West Sumatra and South Borneo, the models typically exhibit poorer performance. Specifically, for Llama–3, the performance gap compared to Bali ranges from 10 to 30%. This highlights the presence of cultural biases and a lack of inclusivity in model reasoning abilities, likely stemming from the distribution of training data. The proximity of West Java to Jakarta (Indonesia’s capital) and Bali’s global status as a tourism destination may contribute to the abundance of textual data on these two cultures.
@@
Table 5 shows the accuracy of the top 3 performing models across different topics. Similar to Table 4, the models perform worse in province-specific contexts for all topics, with the notable exception of food. For province-specific contexts, GPT–4 excels for the themes of food, religious holidays, and arts while for non-specific contexts, Llama–3 achieves accuracy more than 90% for agriculture, daily activities, and religious holidays.
Topic Merak Llama–3 GPT–4
@@
Food 58.7 50.2 69.6 71.8 73.3 77.3
Wedding 60.0 50.2 80.0 66.9 82.5 66.9
Family relationship 55.1 39.6 75.5 66.9 85.2 63.0
Pregnancy and kids 57.1 44.3 83.5 69.0 84.7 64.6
Death 56.9 51.0 86.2 62.5 79.4 63.5
Religious holidays 66.7 64.3 79.7 78.6 91.5 78.6
Agriculture 65.5 54.5 100.0 80.3 93.8 72.7
Farm and fishery 67.9 42.9 85.7 62.3 80.0 72.7
Art 77.8 55.2 83.3 72.2 90.0 82.5
Traditional game 38.5 48.8 69.2 68.3 78.6 64.6
Daily activities 68.5 52.3 92.6 72.7 93.1 76.1
Socio-religious 70.5 51.0 88.5 71.8 86.4 75.8
Table 5: Top-3 model accuracy by topic. “PS” indicates instances containing province-specific contexts, while “$\neg$¬PS” indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
Results by fine-grained cultural elements
We tasked two expert workers with annotating 200 random samples from IndoCulture based on six cultural elements, as derived from Axtell and Fornwald (1998); Williams (2014). [^3] https://www.languageeducatorsassemble.com/elements-of-culture [^4] https://pressbooks.howardcc.edu/soci101/chapter/3-2-the-elements-of-culture/ While these elements may not encompass every cultural aspect, we contend that they cover the most prominent or pivotal elements, including: (1) symbols (material or non-material objects representing meaning); (2) artifacts (material or non-material objects produced by society); (3) values and beliefs (principles, ideas, and concepts assumed to be ideal and correct in society); (4) norms (rules guiding values and beliefs); (5) language; (6) rituals (established procedures and ceremonies); and (7) other, for examples that do not fit into any of the defined elements. This annotation is a multi-label task, and the average Kappa score across the cultural elements is 0.56, with each ranging from 0.4 to 0.75. These scores indicate moderate to substantial agreement.
Table 6 displays the distribution of each cultural element in our dataset, along with the performance breakdown across Merak, Llama–3, and GPT–4. Among the 200 random samples, we observe that 42.5% of our data contains artifacts, 37.5% norms, and 30% rituals. Only 4% of the data pertains to symbols, while 7.5% belongs to the other category. Merak shows lower accuracy in norms, with a 24% decrease compared to values and beliefs. Conversely, Llama–3 performs best in values and beliefs with 85% accuracy, but accuracy drops by 23% for norms. GPT–4 maintains relatively stable accuracies across cultural elements, with differences averaging between 3% and 5%. Furthermore, language presents a challenge for Merak, achieving only 38% accuracy, whereas Llama–3 and GPT–4 achieve 66% and 72% accuracy, respectively.
Cultural element (%) Merak Llama–3 GPT–4
Symbols (4) 50.0 50.0 70.8
Artifacts (42.5) 55.3 74.1 67.8
Values and Beliefs (10.5) 61.9 85.7 69.8
Norms (37.5) 38.7 62.7 73.6
Language (19.5) 38.5 66.7 72.0
Ritual (30) 53.3 65.0 70.7
Other (7.5) 66.7 66.7 69.2
Table 6: Accuracy comparison of Merak, Llama–3–Instruct (70B), and GPT–4 across 200 random samples, categorized by cultural elements. The numerical value following each cultural element indicates its proportion within the samples.
Can the model provide a reasonable explanation to support the answer?
Figure 4: Performance comparison between Merak (7B), Llama–3 Instruct (70B), and GPT–4 based on text generation output. “Answer (T)” indicates that the generated answer is true, while “Exp(F)” denotes that the answer explanation is false.
We conduct a manual investigation of the text generation output for Merak (7B), Llama–3, and GPT–4 across 200 random samples. This involves manually examining the generated answer along with its explanation. To obtain the explanation, we modify the Indonesian prompt in Figure 3 by adding the string Jelaskan jawabanmu! “Explain your answer!”. Our annotation process is binary, categorizing explanations as either True or False. We label an explanation as False if it is absent, contains hallucinations, or provides inaccurate information. [^5] We use the Google search engine to verify the correctness of the explanation.
As anticipated, there is a substantial drop in accuracy for Merak (7B) from 53.2% (as shown in Table 3) to 29.5%. This discrepancy underscores the limitations of relying solely on token probabilities to assess the true capability of a language model. Interestingly, only 4.5% of the samples are answered correctly with the appropriate explanation by Merak, despite it being the top performer among the Indonesian-centric language models. Larger models like Llama–3 and GPT–4 achieve more robust accuracies of 70.5% and 69.5%, which are 3–5% lower than those indicated in Table 3. However, both models encounter challenges in generating appropriate explanations for correctly-answered samples, with Llama–3 and GPT–4 producing explanation errors in 29% and 34% of cases, respectively.
@@
We automatically translated IndoCulture to English using the Google Translate API [^6] Accessed in March 2024. and used the English prompt in Figure 3 to evaluate the models. Specifically for this part, we include more models for comparison. All results over English text dropped, except for Llama–2 and Merak. This could be attributed to two reasons. First, Llama–2 is an English-centric model, and Merak is fine-tuned from Llama–2. Second, the performance drop for other models could be caused by translation errors. We further investigated this with 100 random samples and found that 81 samples had acceptable translations. We observed translation errors such as pronoun mismatches, inaccurate proverb translations, and inaccurate translations of local terms, such as pupuik translated as “fertilizer”.777pupuik is a traditional musical instrument in West Sumatra. It is worth noting that the word pupuik closely resembles the word pupuk in Indonesian, which means fertilizer. To better understand the cultural gap in language models, we followed the approach of Liu et al. (2024) to manually correct the translations and reevaluate the models. We found that GPT–4’s performance over the 100 random samples was 77.0 for the original Indonesian text, 68.0 for the English machine-translated text, and 72.0 for the English translation fixed by humans.